# Curve-Aware Gaussian Splatting for 3D Parametric Curve Reconstruction

### v2

@@ 4.1 Experiment Settings
Datasets. We evaluate the method on three public benchmarks: ABC-NEF [35], MV2Cyl’s real objects [8], and Replica [26]. Consistent with EMAP and EdgeGaussians [5], we utilize a selected subset of the ABC-NEF dataset, comprising 82 CAD models from the ABC dataset [10]. Each model provides parametric ground-truth edges for quantitative 3D edge evaluation, along with 50 rendered views that capture diverse object perspectives. MV2Cyl [8] introduces a collection of real-world objects created by 3D-printing various sketch-extrude models from Fusion360 [30], providing multi-view images, camera poses, and CAD models. We demonstrate 6 objects to evaluate the performance of real-world objects. For Replica, each scene offers 100 multi-view edge maps for a thorough evaluation. Given the lack of ground-truth edge annotations, the analysis on Replica focuses primarily on qualitative comparisons to assess the effectiveness for complex scenes.
@@
Implementation Details. Our method is trained for 10k iterations for each test instance. To ensure meaningful updates, we delay the introduction of adaptive control strategies for parametric edges until later stages of training, as these edges require sufficient optimization to converge to a suitable form. Specifically, curve linearization is initiated at the 3,000th iteration. From the 7,000th iteration onward, we apply curve linearization and merging strategies to refine the parametric curves. To balance efficiency, operations of splitting, linearization, and merging are performed every 1K iterations. Additionally, we optimize the curve opacity before the 7,000th iteration. After that, the opacity attribute is fixed, and the mask loss is applied.
@@ 4.2 Evaluations
Method Runtime $\downarrow$ $N_{\text{Bézier}}$downarrow $N_{\text{line}}$$\downarrow$ $N_{\text{sum}}$$\downarrow$
@@
Table 2: Runtime and number of curves/lines comparisons on ABC-NEF [35]. Results from NEF’s released pretrained models are indicated by $\dagger$. $N_{\text{Bézier}}$, $N_{\text{line}}$ and $N_{\text{curve}}$ represent the average number of reconstructed Bézier, lines, and total curve edges.
@@
Evaluation on ABC-NEF Dataset [35]. The qualitative and quantitative comparisons are presented in Fig. 4 and Tab. 1, respectively. Following EMAP, we quantitatively report the results using PiDiNet [27] and DexiNed [22] as edge detectors, while the visualized results demonstrate the performance of DexiNed. LIMAP achieves impressive precision but lacks the ability to represent curves, often resulting in significant edge omissions when modeling such geometric elements. NEF exhibits lower accuracy due to limitations in point sampling and fitting-based post-processing. EMAP struggles with incomplete reconstructions, resulting in lower scores in completeness and recall. EdgeGaussians frequently produces fragmented or missing reconstructions, primarily due to inaccuracies in estimating principal axes of constructed Gaussians, which cause the subsequent fitting process to converge to suboptimal local minima. Our method directly optimizes parametric edge representations, inherently incorporating local edge constraints and eliminating errors introduced by post-processing fitting stages. As a result, the proposed method outperforms all learning-based methods across all metrics, while achieving precision on par with LIMAP.
To further compare the efficiency and compactness of parametric edge extractionwe also evaluate on the ABC-NEF benchmark, measuring runtime and the number of parametric edges. DexiNed is used as the default image edge detector. Both EdgeGaussians and our method significantly outperform implicit-representation-based approaches like NEF and EMAP in terms of computational efficiency. NEF utilizes the fewest edges due to its frequent edge omissions, as reflected in its low recall scores. Our method achieves compact parametric edge representation with efficient reconstruction while delivering complete and highly accurate edge reconstruction. It shows a superior balance of efficiency, compactness, and reconstruction quality. More comparisons are provided in Supplementary.
Evaluation on MV2Cyl’s real objects  [8]. Real-world datasets often pose significant challenges to algorithms due to inconsistencies in edge detection across different viewpoints caused by varying perspectives, as well as difficulties in ensuring uniform coverage of objects within the camera’s field of view. Since NEF and EMAP fail to produce reasonable 3D edges in our experiments, we compare our approach with EdgeGaussian. As shown in Fig. 1, our method demonstrates better reverse engineering capabilities. EdgeGaussian generates a significant amount of noisy Gaussian distributions due to the introduction of 2D edge noise caused by variations in lighting and viewpoint in real-world scenes. In contrast, our method inherently enforces local consistency constraints and better mitigates the impact of noise. For qualitative evaluation, we follow MV2Cyl [8] and use ICP to align the reconstructed edge point cloud with the ground-truth CAD model. Tab. 3 shows that our method achieves higher accuracy and more compact 3D edges.
@@ 5 Conclusion
This paper presents a one-stage 3D parametric curve reconstruction method based on a bi-directional coupling mechanism between parametric curves and edge-oriented Gaussians. The limitation is that the method remains contingent on the quality of 2D edge maps, and it may be mitigated through 2D foundation models for estimating better edge maps. Our method also provides a potential pathway and future direction for lifting 2D edge detection foundation models into 3D, to benefit 3D visual and geometric reasoning.

### v3

@@ 4.1 Experiment Settings
Datasets. We evaluate the method on three public benchmarks: ABC-NEF [52], MV2Cyl’s real objects [14], and Replica [34]. Consistent with EMAP [19] and EdgeGaussians [6], we utilize a selected subset of the ABC-NEF dataset, comprising 82 ABC [16] models. Each model provides parametric ground-truth edges for quantitative 3D edge evaluation, along with 50 rendered views that capture diverse object perspectives. MV2Cyl [14] introduces a collection of real-world objects created by 3D-printing various sketch-extrude models from Fusion360 [43], providing multi-view images, camera poses, and CAD models. We demonstrate 6 objects to evaluate the performance of real-world objects. For Replica, each scene offers 100 multi-view edge maps for a thorough evaluation. Given the lack of ground-truth edge annotations, the analysis on Replica focuses primarily on qualitative comparisons to assess the effectiveness for complex scenes.
@@
Implementation Details. Our method is trained for 10k iterations for each test instance. To ensure meaningful updates, we delay the introduction of adaptive control strategies for parametric edges until later stages of training, as these edges require sufficient optimization to converge to a suitable form. Specifically, curve linearization is initiated at the 3k-th iteration. From the 7k-th iteration onward, we apply curve linearization and merging strategies to refine the parametric curves. To balance efficiency, operations of splitting, linearization, and merging are performed every 1k iterations. Additionally, we optimize the curve opacity before the 7k-th iteration. After that, the opacity attribute is fixed, and the mask loss is applied.
@@ 4.2 Evaluations
[…]
@@
[…]
@@
Evaluation on ABC-NEF Dataset [52]. The qualitative and quantitative comparisons are presented in Fig. 4 and Tab. 1, respectively. Following EMAP [19], we quantitatively report the results using PiDiNet [35] and DexiNed [29] as edge detectors, while the visualized results demonstrate the performance of DexiNed. LIMAP [22] achieves impressive precision but lacks the ability to represent curves, often resulting in significant edge omissions. NEF [52] exhibits lower accuracy due to limitations in point sampling and fitting-based post-processing. EMAP struggles with incomplete reconstructions, resulting in lower scores in completeness and recall. EdgeGaussians [6] frequently produces fragmented reconstructions, due to inaccuracies in estimating principal axes of constructed Gaussians, which cause the subsequent fitting to converge to suboptimal local minima. Our method directly optimizes parametric edge representations, inherently incorporating local edge constraints and eliminating errors introduced by fitting stages. As a result, the proposed method outperforms all learning-based methods across all metrics, while achieving precision on par with LIMAP.
To further compare the efficiency and compactness of parametric edge extraction, we also evaluate runtime and the number of parametric edges using DexiNed as the image edge detector. Both EdgeGaussians and our method significantly outperform implicit-representation-based approaches like NEF and EMAP in terms of computational efficiency. NEF utilizes the fewest edges due to its frequent edge omissions, as reflected in its low recall scores. Our method achieves compact parametric edge representation with efficient reconstruction while delivering complete and highly accurate edge reconstruction. It shows a superior balance of efficiency, compactness, and accuracy.
Evaluation on MV2Cyl’s real objects  [14]. Real-world datasets often pose significant challenges to algorithms due to inconsistencies in edge detection across different viewpoints and non-uniform coverage of objects. Since NEF and EMAP fail to produce reasonable 3D edges in our experiments, we compare our approach with EdgeGaussian. As shown in Fig. 1, our method demonstrates better reverse engineering capabilities. EdgeGaussian generates a significant amount of noisy Gaussian distributions due to the introduction of 2D edge noise caused by variations in lighting and viewpoint in real-world scenes. In contrast, our method inherently enforces local consistency constraints and better mitigates the impact of noise. We follow MV2Cyl [14] and use ICP to align the reconstructed edge point cloud with the ground-truth CAD model. Tab. 3 shows that our method achieves higher accuracy and more compact 3D edges.
@@ 5 Conclusion
This paper presents a one-stage 3D parametric curve reconstruction method based on a bi-directional coupling mechanism between parametric curves and edge-oriented Gaussians. The limitation is that the method remains contingent on the quality of 2D edges, and it may be mitigated through 2D foundation models for estimating better edge maps. Our method also provides a potential pathway and future direction for lifting 2D edge detection foundation models into 3D, to benefit 3D visual and geometric reasoning.
Acknowledgements
This work is supported in part by the NSFC (62325211, 62132021, 62372457), the Major Program of Xiangjiang Laboratory (23XJ01009), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Natural Science Foundation of Hunan Province of China (2022RC1104).
Supplementary Material
This supplementary material first provides detailed experimental settings, including data processing procedures, implementation details for comparison baseline methods, and evaluation metrics.
In Fig. 7, we provide additional visual comparisons of our method against state-of-the-art baselines on ABC-NEF [52]. Our code and data are available at https://github.com/zhirui-gao/Curve-Gaussian.
6 Datasets
The proposed method is evaluated on three publicly available datasets: ABC-NEF [52], Mv2Cyl’s Real Objects [14], and the Replica Dataset [34]. Detailed descriptions of each dataset and the experimental setups are provided below.
ABC-NEF Dataset. The ABC-NEF dataset is a widely adopted benchmark for evaluating 3D curve reconstruction quality. It contains precise CAD models with diverse curve types, comprising 115 objects in total. Following the protocols of EMAP [19] and EdgeGaussians [6], objects with indistinct sharp curve features were filtered out, resulting in 82 objects for evaluation. Each object includes 50 images with a resolution of $800\times 800$. For computational efficiency, all images were resized to $400\times 400$.
Mv2Cyl’s Real Objects Dataset. The Mv2Cyl’s Real Objects dataset consists of multiple 3D-printed objects captured using an iPhone 12. Multi-view images were extracted from continuous videos, with camera poses computed using COLMAP. Ground truth CAD models are also provided for evaluation. This dataset presents a significant challenge for 3D edge detection due to the top-down perspective of the captured images. To generate edge maps, the following pipeline is employed:
  • SAM2 [30] is utilized to segment objects from the background.
  • A monocular normal estimation network [50] is applied to identify high-curvature regions as object edges. This approach demonstrated superior performance compared to edge detection methods [35, 29], which often extracted irrelevant edges due to lighting interference.
  • Edge maps are resized to $480\times 480$ for faster processing.
Since the camera poses and CAD models are not aligned in the same coordinate system, the Iterative Closest Point (ICP) was used to register the reconstructed edge points with the ground truth CAD curves for quantitative evaluation.
The lack of real-world benchmarks for 3D curve reconstruction is a notable gap in the field. The Mv2Cyl’s Real Objects dataset addresses this limitation. After getting permission from the authors of Mv2Cyl, a standardized benchmark for evaluating 3D edge reconstruction will be proposed based on their multi-view images. We believe it will be an important contribution to the field.
Replica Dataset. We follow the experimental setup in EMAP [19] on this dataset, focusing on three scenes: Room 0, Room 1, and Room 2.
7 Baselines
Our method is compared against four state-of-the-art 3D line and curve reconstruction baselines. These include three learning-based methods—NEF [52], EdgeGaussians [6], and EMAP [19]—and one line-based Structure-from-Motion (SfM) method, LIMAP [22]. For a fair comparison, the default parameters and settings provided by the authors are adopted for all baselines. In the case of NEF and EMAP, their pre-trained models were directly applied to generate visual results.
8 Evaluation Metrics
To quantitatively assess the performance of our method, we adopt a set of evaluation metrics that align with established protocols in this field. Points are uniformly sampled along both the reconstructed parametric curves and the corresponding ground-truth edges, enabling a direct comparison between them. The metrics are introduced as follows:
  • Accuracy: This metric calculates the average distance from each predicted point to its closest counterpart on the ground-truth curve. Smaller values correspond to higher accuracy.
  • Completeness: This measures the average distance from each ground-truth point to the nearest predicted point. Improved performance is indicated by lower values.
  • Precision at Threshold $\tau$ ($P(\tau)$): This quantifies the proportion of predicted points that lie within a distance $\tau$ of any ground-truth point. Higher precision values reflect better alignment with the ground truth.
  • Recall at Threshold $\tau$ ($R(\tau)$): This evaluates the proportion of ground-truth points that have at least one predicted point within a distance $\tau$. Higher recall values signify better coverage of the ground truth.
In addition to these conventional metrics, we introduce a new metric, Curve Count, which evaluates the compactness of the reconstructed curves, which can be found in Table 2 of the main paper. This metric counts the total number of curves generated by the method, providing insight into the efficiency of the representation. A smaller curve count indicates a more concise and compact reconstruction, which is advantageous for downstream tasks that require efficient curve processing.
For consistency with prior evaluations, precision and recall are computed at distance thresholds of $\tau=5$, $10$, and $20$ millimeters (mm). For other parameters in the evaluation, such as the number of sampling points, we adhere to the same settings as those used in EMAP [19].
9 Implementation Details
The weight coefficients $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ were set to $0.01$, $0.01$, $0.01$, and $0.0005$, respectively. For all Bézier curves and straight lines, a default of 12 Gaussian points was sampled per curve.
For the ABC dataset, the midpoints of Bézier curves were initialized by uniformly sampling $15\times 15\times 15$ points in 3D space. In contrast, for COLMAP-based datasets, the midpoints were initialized using the point cloud generated by Structure-from-Motion (SfM). The threshold for curve merging was kept consistent with EMAP [19].
To ensure the quality of the reconstructed curves, curves with an opacity below $0.05$ were removed, and curves with bending angles exceeding $20^{\circ}$ were split. Additionally, Gaussian components with mask attributes below $0.01$ were considered redundant and discarded. Further details on parameter design and implementation can be found in the accompanying code.
Figure 7: Qualitative comparisons on ABC-NEF [52]. Distinct colors represent different curves. Our method achieves more complete and accurate edge reconstruction of objects while maintaining parametric compactness.